# New Method for Robust Analysis of Critical Behavior of Magnetic Systems

### v1

@@ 2. Modified Arrott plots
where $M_{1}$ is a temperature and field-dependent constant. $T$ is the temperature at which the M-H isotherms are recorded. The plots of $(\mu_{0}H/M)^{1/{\gamma}}$ vs $M^{1/{\beta}}$ is known as modified Arrott plots (MAPs).
@@ 3. Iteration method
The concept to develop the empirical relation 1 was to yield quasi straight line in the saturation magnetization region using a particular set of $\beta$ and $\gamma$ in the vicinity of $T_{\rm{C}}$. To get appropriate $\beta$ and $\gamma$, an iteration process was employed as follows:
@@
respectively. Here, $\epsilon=\displaystyle\frac{T-T_{\rm{C}}}{T_{\rm{C}}}$ is the reduced temperature, $M_{0}$ and $h_{0}$ are constants. The above fittings (Eqs. 4 and 5) will yield new values of $\beta$ and $\gamma$. Again, a new MAPs are constructed using the obtained new values of $\beta$ and $\gamma$. Further, following the above processes (Eqs. 2 to 5), another set of $\beta$ and $\gamma$ are generated. The above processes will be repeated till the converged values of $\beta$ and $\gamma$ are achieved.
@@ 4. Modified iteration method
Modified iteration method (MIM) [15] yields two sets of $\beta$ and $\gamma$ on either side of $T_{\rm{C}}$. The process to estimate appropriate $\beta$ and $\gamma$ using MIM is as follows.
(i) First, including the M-H isotherm taken at $T_{\rm{C}}$, make two sets of M-H isotherms: one for $T\leq T_{\rm{C}}$ and another one for $T\geq T_{\rm{C}}$.
(ii) Take one set of M-H isotherms. Now, using $\beta$ and $\gamma$ of closest standard UC, estimate normalized slopes (NSs) as discussed in the main text. One should keep in mind--vary $\beta$ and $\gamma$ in such a way that the intercept, estimated from the linear fit to the MAPs at $T_{\rm{C}}$, should pass through origin.
(iii) Take another values of $\beta$ and $\gamma$ (either by increasing or decreasing) in such a way that constructed NSs get closer to 1.
(iv) Stop the process when NSs become closest to 1 within error limit. This is how one set of $\beta$ and $\gamma$ will be estimated.
(v) Repeat steps (ii) to (iv) for the remaining set of M-H isotherms.
Thus, two separate sets of $\beta$ and $\gamma$ will be estimated.
@@ 5. New Proposals
a. The iterative process
In the main text, we propose a new method to get appropriate CEs of a system under investigation. The proposal correlates $\beta$ and corresponding $\gamma$ with $\delta$ as follows
@@
where $\delta$ is estimated from the M-H at $T_{\rm{C}}$ in different field ranges using the relation
@@
where $M_{0}$ is a proportionality constant. Thus, the corresponding $\gamma$ is obtained by putting $\beta$ for a fix $\delta$, which is determined from the M-H at $T_{\rm{C}}$, in Eq. 6. In this way, one can get a series of $\beta$ and corresponding $\gamma$. Here, we have varied $\beta$ with iteration number, n as $\beta(n)=\beta_{i}+0.001\times n$. In this way, one gets error in the value of $\beta$ in third decimal place. The multiplication factor, $0.001$ is depend on the choice of the user. Thus, we control the error in $\beta$ manually. The error in $\gamma$ depends on imposed error in $\beta$ and the error appearing in $\delta$. It is obvious that one may get error in the magnetization during experiment. For the same, we put error in $\beta$ up to $0.003$. The reverse calculation using relation 6 is also true, i.e., vary $\gamma$ and determine corresponding $\beta$ as $\beta=\displaystyle\frac{\gamma}{\delta-1}$. In this way, one can control the error in $\gamma$ manually and the error in $\beta$ will depend on the imposed error in $\gamma$ and the error appearing in $\delta$.
Finally, we construct the MAPs using the final CEs and by selecting the temperature range $T\leq T_{\rm{C}}$. The linear fit to MAPs in the selected field-range yields slopes and intercepts. As observed and has been discussed in the main text, intercepts make crossover at $T_{\rm{C}}$, which means, linear fit to MAP at $T_{\rm{C}}$ will pass through origin. After estimating $NS$, which is defined as $NS=\displaystyle\frac{\text{slope of the linear fit to MAP at }T}{\text{slope of the linear fit to MAP at }T_{\rm{C}}}$, we determine average of NS ($NS_{\rm{avg}}$) as $NS_{\rm{avg}}=\frac{\sum_{i=1}^{N}(NS)_{i}}{N}$, where $i$ denotes the $M-H$ isotherm at a given $T$, and $N$ is the total number of the isotherms appearing in the calculation. As shown in Fig. S1, those values of $\beta$ and corresponding $\gamma$ (Eq. 6) are the appropriate for which $NS_{\rm{avg}}\sim 1$. The details are mentioned in the main text.
@@
Moreover, we have related $\beta$ with $n$ as $\beta(n)=\beta_{i}+0.001\times n$, where $\beta_{i}=0.400$ [Fig. S1(a)] for the selected field range. One can get appropriate $n$ by putting $NS_{\rm{avg}}=1$, slope$=-0.002$ and intercept$=1.016$ in Eq. 8. Thus, from the linear fit (Eq. 8), we get $n=8$ using which we determine $\beta=0.408$ which matches exactly as obtained from the conventional process.
In the unconventional approach, the iteration is not required, instead one can select a few set of $\beta$ and $\gamma$ correlated with Eq. 6, in such a way that the estimated $NS_{\rm{avg}}$ falls between 0.95 and 1.05. One can go for the selection of the $NS_{\rm{avg}}$ range from 0.9 to 1.1 or more, but it should be noted that the $NS_{\rm{avg}}$ vs. $\beta$ plot should show quasi-linear variation in the vicinity of $NS_{\rm{avg}}=1$ as shown in Fig. S2 and Fig. 1(h) in the main text. Now, estimate $NS_{\rm{avg}}$ for the selected set of $\beta$ and corresponding $\gamma$, and do $NS_{\rm{avg}}$ vs. $\beta$ plot. Here, one can take a minimum sets of $\beta$ and corresponding $\gamma$ to perform a linear fit. Our, recommendation is to select at least three set of $\beta$ and corresponding $\gamma$ in such a way the estimated $NS_{\rm{avg}}$ falls on both side of $NS_{\rm{avg}}=1$. Finally, the linear fit to the plot will yield $slope$ and $intercept$. Now, using the relations,
@@
the appropriate respective values of $\beta$ and $\gamma$ can be determined easily. We have tested this concept for the field range 10 to 40 T as shown in Fig. S2. As shown, from the intersection point of the linear fit to data and the horizontal line drawn passing through $NS_{\rm{avg}}=1$, we directly get $\beta=0.3435$. By inserting the determined $slope$ and $intercept$ (Fig. S2) in relation 9, we get $\beta\approx 0.343$, which matches exactly as determined using our proposed conventional iteration process and SP. However, when the range of $NS_{\rm{avg}}$ is increased up to $1.00\pm 0.15$ [Inset of Fig. S2], the determined value of $\beta\approx 0.347$ starts deviating away. And $NS_{\rm{avg}}$ vs. $\beta$ started showing curvature.
@@ 6. Normalized slope plots for Ni
Figure S3 shows the normalized slope (NS) vs. temperature plots determined from the slope of the linear fit the modified Arrott-plots constructed using the CEs of standard UCs and the CEs determined using standard procedure (SP).
@@ 7. Field-dependent normalized slope and intercept plots for SC Gd
The linear fits to MAPs, constructed using the CEs of standard UCs and the CEs determined using our new proposal (NP) [Fig. S1], yield slopes and intercepts. Figure S4 represents the variation of NSs with temperature (defined above and in the main text) estimated using the CEs of standard UCs and the CEs determined using NP. As shown in Fig. S4, NSs make the first indication that the critical behavior of SC should belong to either 3D XY UC or 3D Heisenberg UC in low-field range. But, the determined CEs differ the guess made by NS plots. The primary guess made by the NS plots for higher field range [Figs. S4(b) to S4(f)], which is 3D Ising UC, do not depict the full physics of the system. The NS plots, constructed using the CEs determined using NP, are always closest to 1, which is the indication of the appropriate CEs. The determined CEs [Fig. S1] do not belong to a single UC. This is the indication that one cannot depict the emerging physics directly from the NS plots using the CEs of standard UCs. One need to perform the critical analysis to reveal the correct physics.
@@
Figure S5 represents the variation of Intercepts with temperature (defined above and in the main text) estimated using the CEs of standard UCs and the CEs determined using NP. Let $T_{\rm{1}}$ is the temperature where the intercepts for short-range UCs (3D Heisenberg, 3D Ising and 3D XY) make crossover and has very less variation with applied field range. Let $T_{\rm{2}}$ is the temperature where the intercepts for long-range MF UC make crossover and has significant variation with applied field range. We have corroborated in the main text that the variation of $T_{\rm{2}}$ indicate the dominance of long-range coupling in low-field range which eventually gets weaker with applied field resulting in the coexistence of both long- and short-range coupling in the high-field range. The intercepts, estimated using the CEs determined using NP, make crossover at $T_{\rm{C}}$.
@@ 8. Modified Arrott plots: theoretical and experimental results
Figure S6 represents the MAPs constructed for Ni using the CEs of standard UCs [Figs. S6(a) to S6(d)] and the CEs determined by standard process (SP) [Fig. S6(e)], iteration method (IM) [Fig. S6(f)], MIM [Fig. S6(g)], and NP [Fig. S6(h)] for the critical analysis. As one can see the MAPs constructed using the CEs determined from IM show slightly non-co-linear variation which is also evidenced from the NS plots [Fig. S3(b)] which show deviation from 1. While the MAPs constructed using the CEs determined from MIM and NP show quasi linear variation for all the isotherms for below $T_{\rm{C}}$. This is also evidenced from the NS plots [Figs. S3(c) and S3(d)] which is close to 1 for below $T_{\rm{C}}$. As observed, the linear variation of MAPs, constructed using the CEs determined from MIM and NP, for above $T_{\rm{C}}$ show that the field-range of the linear variation decreases with increasing temperature. This investigation yields two points: (i) one should do the critical analysis in low-temperature and high-field ranges for above $T_{\rm{C}}$, and (ii) to perform critical analysis in the whole temperature range, the field-range should be varied to acquire the linear variation in the MAPs.
@@
Figure S7 represents the MAPs constructed for SC Gd system using the CEs of standard UCs [Figs. S7(a) to S7(e)] and the CEs determined by the NP for different field ranges [Figs. S7(f) to S7(j)] for below $T_{\rm{C}}$. The MAPs, constructed using the CEs of standard UCs, indicate that the critical behavior of SC Gd should belong to one of the short-range standard UCs. The MAPs, constructed using the CEs determined by NP, for the field range 1 to 2 T show significant but smaller deviation from quasi-linear variation. The MAPs for higher field-ranges show almost quasi-linear variation. This is the indication that the results obtained for the field range 1 to 2 T may be ambiguous. The critical analysis for above $T_{\rm{C}}$ have been ignored due to the reasons discussed in the main text while performing theoretical investigation on Ni system.
@@ 9. Scaling results
where $\tilde{h}(x)$ is the scaling function. Thus, the relation 11 stattes that for appropriate CEs and $T_{\rm{C}}$, the plots of $M(\mu_{0}H)^{1/\delta}$ vs. $\epsilon$$(\mu_{0}$)${}^{-\frac{1}{\beta+\gamma}}$ should corresponds to a universal curve onto which all the isotherms collapse. The Eq. 11 may be formally converted as
@@
where $f_{+}$ and $f_{-}$ are defined for $T>T_{\rm{C}}$ and $T<T_{\rm{C}}$, respectively. The relation 12 states -- for appropriate values of $\beta$ and $\gamma$, the plots of renormalized magnetization, $m=M(H,\epsilon)|\epsilon|^{-\beta}$ vs. renormalized field, $h=\mu_{0}H|\epsilon|^{-(\beta+\gamma)}$ should correspond to collapse of the isotherms on two separate universal curves for below and above $T_{\rm{C}}$. Equation 12 is most commonly used relation to verify the appropriateness of the obtained CEs. It has been proven [15] that these scaling relations (Eqs. 11 and 12) can absorb significantly large errors appearing in the determination of CEs. So, a new and appropriate relation is needed to check the reliability of determined CEs. However, Widom relation, $\delta=1+\displaystyle\frac{\gamma}{\beta}$, may be used as a primary tool to check the reliability and accuracy of determined CEs using various methods.
@@
Figures S8(a) to S8(d) show the plots of $m$ vs. $h$ constructed using the CEs determined by SP, IM, MIM and NP for Ni system. As one can see clearly that the $m$ vs. $h$ plots using the CEs determined by IM [Fig. S8(b)] show maximal deviation from overlapping on two separate isotherms. The $m$ vs. $h$ using the CEs determined by SP [Fig. S8(a)] have perfect collapse on two universal curves for the whole range of the M-H isotherms. Similar perfect collapse on two universal curves for the whole range of the M-H isotherms have been observed using the CEs determined by MIM [Fig. S8(c)] and NP [Fig. S8(d)] for below $T_{\rm{C}}$. We have used the same CEs, determined by MIM and NP for below $T_{\rm{C}}$, to construct the $m$ vs. $h$ plots for above $T_{\rm{C}}$, which show slight deviation from collapsing on a universal curve. Figures S8(e) to S8(h) are the renormalized Arrott-plots (RAPs) constructed using the CEs determined by SP, IM, MIM and NP. The perfect overlapping is closely observable for the CEs determined by SP, and MIM and NP for below $T_{\rm{C}}$. The RAPs constructed using the CEs determined by IM, MIM for above $T_{\rm{C}}$ and NP for above $T_{\rm{C}}$ show slightly large deviation from collapsing on universal curve. These are the indication that the CEs determined using SP, and MIM and NP for below $T_{\rm{C}}$ are appropriate to reveal the properties of a system under investigation. In fact, the CEs determined from the above three mentioned methods (SP, MIM and NP) match significantly to each other within error limit. The details have been discussed in the main text. However, no concrete conclusion can be made from the use of relation 11 because the constructed plots using the CEs determined by all methods overlap significantly onto a single universal curves as shown in Figs. S8(i) to S8(l). From these investigations, One important information is coming out -- the overlapping of all the isotherms on universal curves for entire field-range is the indication of the existence of single exchange coupling in the system, which was intentionally incorporated during simulation.
@@
Figure S9 represents the scaling plots and RAPs for SC Gd. One should note that the critical analysis on SC Gd have been carried out only for below $T_{\rm{C}}$ but we have plotted the above $T_{\rm{C}}$ isotherms also using the same CEs as determined for below $T_{\rm{C}}$. The plots shown in Fig. S9, have been constructed using the CEs determined by NP. The $m$ vs. $h$ plots show nice overlapping in the field-ranges which were used to determine CEs for below $T_{\rm{C}}$. However, in the low field range (approximately below 0.5 T) the $m$ vs. $h$ plots [Figs. S9(a) to S9(f)] show deviations from collapsing on the universal curves. In the similar fashion, the RAPs [Figs. S9(g) to S9(l)] also sow overlapping in the high field range while deviations are observable in the low field range. The scaling plots [Figs. S9(m) to S9(r)], constructed using the relation 11, also show significant deviations in the low-field range mainly below $T_{\rm{C}}$. The reported scaling results using Eq. 11 do not show deviation from collapsing on single universal curve. The reason behind this could be the skipping of the low-field data. However, from the deviation in the low-field region, we state that this is the signature of the competing energies along with symmetric exchange interaction in SC Gd. The detail discussions have been given in the main text.

### v2

@@ 2. Modified Arrott plots
where $M_{1}$ is a temperature and field-dependent constant. $T$ is the temperature at which the M-H isotherms are recorded. The plots of $(\mu_{0}H/M)^{1/{\gamma}}$ vs $M^{1/{\beta}}$ is known as modified Arrott plots (MAPs), see below Figs. S6 and S7.
@@ 3. Iteration method
The concept to develop the empirical relation 1 was to yield quasi linear behavior of MAPs [Figs. S6 and S7] in the saturation magnetization region using a particular set of $\beta$ and $\gamma$ in the vicinity of $T_{\rm{C}}$. Later, an IM was employed to determine appropriate values of the CEs. To get appropriate $\beta$ and $\gamma$, an iteration process was employed as follows:
@@
respectively. Here, $\epsilon=\displaystyle\frac{T-T_{\rm{C}}}{T_{\rm{C}}}$ is the reduced temperature, $M_{0}$ and $h_{0}$ are constants. The above fittings (Eqs. 4 and 5) yield new values of $\beta$ and $\gamma$. Again, a new MAPs are constructed using the obtained values of $\beta$ and $\gamma$. Further, following Eqs. 2 to 5, another set of $\beta$ and $\gamma$ are generated. The above processes will be repeated till the converged values of $\beta$ and $\gamma$ are achieved.
@@ 4. Modified iteration method
MIM yields two sets of $\beta$ and $\gamma$ on either side of $T_{\rm{C}}$ [17]. The process to estimate appropriate $\beta$ and $\gamma$ using MIM is as follows.
  1. 1.
    First, including the M-H isotherm taken at $T_{\rm{C}}$, make two sets of M-H isotherms: one for $T\leq T_{\rm{C}}$ and another for $T\geq T_{\rm{C}}$.
  2. 2.
    Take one set of M-H isotherms. Now, using $\beta$ and $\gamma$ of a standard UC, estimate normalized slopes (NSs) as discussed in the main text. One should keep in mind--vary $\beta$ and $\gamma$ in such a way that the intercept, estimated from the linear fit to the MAP at $T_{\rm{C}}$, should pass through origin.
  3. 3.
    Take another values of $\beta$ and $\gamma$ (either by increasing or decreasing) in such a way that constructed NSs get closer to 1.
  4. 4.
    Stop the process when NSs become closest to 1 within error limit. This is how one set of $\beta$ and $\gamma$ will be estimated.
  5. 5.
    Repeat steps (2) to (4) for the remaining set of M-H isotherms.
Thus, two separate sets of $\beta$ and $\gamma$ are estimated. In this article, we claim that the critical behavior remains the same on either side of $T_{\mathrm{C}}$. The discrepancy reported in [17], may have come due to (i) closest values of NS, within error limit, above $T_{\mathrm{C}}$, and (ii) selection of little lower temperature range ($\approx$ 6 K) for above $T_{\mathrm{C}}$ as compared to the selected temperature range of 70 K, here, for Ni during computational investigation.
@@ 5. New Proposals
a. The successive process
In the main text, we propose a new method to get appropriate values of CEs for magnetic systems. The proposal correlates $\beta$ and $\gamma$ with $\delta$ as follows
@@
where $\delta$ is estimated from the M-H at $T_{\rm{C}}$ using the relation
@@
where $M_{0}$ is a proportionality constant. It is necessary to estimate $\delta$ for different field range if one wants to perform field-dependent investigation of critical behavior. Our recommendation is to check if the value of $\delta$ changes with field or not. Then, depending on the situation, one need to perform critical analysis. Then, determine corresponding $\gamma$ by putting $\beta$ and $\delta$, which is determined using relation 7, in Eq. 6. Thus, one can get a series of $\beta$ and corresponding $\gamma$. Here, we have varied $\beta$ with iteration number, n as $\beta(n)=\beta_{i}+0.001\times n$. In this way, one gets error in the value of $\beta$ in third decimal place. The multiplication factor, $0.001$ is dependent on the choice of the user. Thus, we control the error in $\beta$ manually as has been discussed in the main text. It is obvious that one may get error in the magnetization during experiment. For the same, we put error in $\beta$ up to $0.003$. The reverse calculation using relation 6 is also true, i.e., vary $\gamma$ and determine corresponding $\beta$ as $\beta=\displaystyle\frac{\gamma}{\delta-1}$. In this way, one can control the error in $\gamma$ manually and the error in $\beta$ will depend on the imposed error in $\gamma$ and the error appearing in $\delta$.
Finally, we construct the MAPs using the sets of the CE values as discussed above and by selecting the temperature range, $T\leq T_{\rm{C}}$. The linear fit to MAPs in the selected field-range yields slopes and intercepts. As observed and has been discussed in the main text, intercepts make crossover at $T_{\rm{C}}$, which means, linear fit to MAP at $T_{\rm{C}}$ will pass through origin. After estimating $NS$, which is defined as $NS=\displaystyle\frac{\text{slope of the linear fit to MAP at }T}{\text{slope of the linear fit to MAP at }T_{\rm{C}}}$, we determine average of NS ($NS_{\rm{avg}}$) as $NS_{\rm{avg}}=\frac{\sum_{i=1}^{N}(NS)_{i}}{N}$, where $i$ denotes the $M-H$ isotherm at a given $T$, and $N$ is the total number of the isotherms appearing in the calculation. As shown in Fig. S1, those values of $\beta$ and corresponding $\gamma$ (Eq. 6) are appropriate for which $NS_{\rm{avg}}\sim 1$. The details are mentioned in the main text. One can follow the same process for $T\geq T_{\rm{C}}$. However, we have shown in the main text that analysis for $T\geq T_{\rm{C}}$ is not required.
@@
Moreover, we have related $\beta$ with $n$ as $\beta(n)=\beta_{i}+0.001\times n$, where $\beta_{i}=0.400$ [Fig. S1(a)] for the selected field range. One can get appropriate $n$ by putting $NS_{\rm{avg}}=1$, slope$=-0.002$ and intercept$=1.016$ in Eq. 8. Thus, from the linear fit (Eq. 8), we get $n=8$ using which we determine $\beta=0.408$ which matches exactly as obtained using the successive process.
In the unconventional approach, the succession process is not required, instead we select a few set of $\beta$ and corresponding $\gamma$ (correlated with Eq. 6) in such a way that the estimated $NS_{\rm{avg}}$ falls between 0.95 and 1.05. One can go for the selection of the $NS_{\rm{avg}}$ range from 0.9 to 1.1 or more, but it should be noted that the $NS_{\rm{avg}}$ vs. $\beta$ plot should show quasi-linear variation in the vicinity of $NS_{\rm{avg}}=1$ as shown in Fig. S2. Now, estimate $NS_{\rm{avg}}$ for the selected set of $\beta$ and corresponding $\gamma$, and do $NS_{\rm{avg}}$ vs. $\beta$ plot. Here, one can take a minimum sets of $\beta$ and corresponding $\gamma$ to perform a linear fit. Our, recommendation is to select at least three set of $\beta$ and corresponding $\gamma$ in such a way the estimated $NS_{\rm{avg}}$ falls on both side of $NS_{\rm{avg}}=1$. Finally, the linear fit to the plot will yield $slope$ and $intercept$. Now, using the relations,
@@
the appropriate respective values of $\beta$ and $\gamma$ can be determined easily. We have tested this concept for the field range 10 to 40 T as shown in Fig. S2. As shown, from the intersection point of the linear fit to data and the horizontal line drawn passing through $NS_{\rm{avg}}=1$, we directly get $\beta=0.3435$. By inserting the determined $slope$ and $intercept$ (Fig. S2) in relation 9, we get $\beta\approx 0.343$, which matches exactly as determined using our proposed successive process for the estimation of CEs and the SP. However, when the range of $NS_{\rm{avg}}$ is increased up to $1.00\pm 0.15$ [Inset of Fig. S2], the determined value of $\beta\approx 0.347$ starts deviating. And $NS_{\rm{avg}}$ vs. $\beta$ started showing curvature.
@@ 6. Normalized slope plots for Ni
Figure S3 shows the normalized slope (NS) vs. temperature plots obtained using the slope of the MAPs constructed using the CE values of the standard UCs and the values determined using standard procedure (SP). As one can see, Fig. S3 (b), the NSs show significant deviation from $1$ on either side of $T_{\mathrm{C}}$ when obtained using the values of CEs as determined using the IM.
@@ 7. Field-dependent normalized slope and intercept plots for SC Gd
Figure S4 represents the variation of NSs with temperature (defined above and in the main text) estimated using the CEs of standard UCs and the CEs determined using NP. As shown in Fig. S4, NSs make the first indication that the critical behavior of SC should belong to either 3D XY UC or 3D Heisenberg UC in low-field range. But, the determined CEs differ the guess made by NS plots. The primary guess made by the NS plots for higher field range [Figs. S4(b) to S4(f)], which is 3D Ising UC, do not depict the full physics of the system. The NS plots, constructed using the CEs determined using NP, are always closest to 1, which is the indication of the appropriate CEs. The determined CEs [Fig. S1] do not belong to any single UC. This is the indication that one cannot depict the emerging physics directly from the NS plots using the CEs of standard UCs. Therefore, one need to perform the critical analysis to reveal the emerging physics.
@@
Figure S5 represents the variation of Intercepts with temperature (defined above and in the main text) estimated using the CEs of standard UCs and the CEs determined using NP. Let $T_{\rm{1}}$ is the temperature where the intercepts for short-range UCs (3D Heisenberg, 3D Ising and 3D XY) make crossover. The variation of $T_{\rm{1}}$ with applied field range is almost constant. Let $T_{\rm{2}}$ is the temperature where the intercepts for long-range MF UC make crossover. The values of $T_{\rm{2}}$ show significant variation with applied field range. We have corroborated in the main text that the variation of $T_{\rm{2}}$ indicate the dominance of long-range coupling in low-field range which eventually gets weaker with applied field resulting in the coexistence of both long- and short-range coupling in the high-field range. The intercepts, estimated using the CEs determined using NP, make crossover at $T_{\rm{C}}$.
@@ 8. Modified Arrott plots: theoretical and experimental results
Figure S6 represents the MAPs constructed for Ni using the CEs of standard UCs [Figs. S6(a) to S6(d)] and the CEs determined by standard process (SP) [Fig. S6(e)], iteration method (IM) [Fig. S6(f)], MIM [Fig. S6(g)], and NP [Fig. S6(h)] for the critical analysis. As one can see the MAPs constructed using the CEs determined from IM show slightly non-co-linear variation which is also evidenced from the NS plots [Fig. S3(b)] which show deviation from 1. While the MAPs constructed using the CEs determined from MIM and NP show quasi linear variation for all the isotherms for below $T_{\rm{C}}$. This is also evidenced from the NS plots [Figs. S3(c) and S3(d)] which is close to 1 for below $T_{\rm{C}}$. As observed, the linear variation of MAPs, constructed using the CEs determined from MIM and NP, for above $T_{\rm{C}}$ indicate the field-range of the linear variation decreases with increasing temperature. This investigation yields two points: (i) one should do the critical analysis in low-temperature and high-field ranges for above $T_{\rm{C}}$, and (ii) to perform critical analysis in the whole temperature range, the field-range should be varied to acquire the linear variation in the MAPs.
@@
Figure S7 represents the MAPs constructed for SC Gd using the CEs of standard UCs [Figs. S7(a) to S7(e)] and the CEs determined by the NP for different field ranges [Figs. S7(f) to S7(j)] for below $T_{\rm{C}}$. The MAPs, constructed using the CEs of standard UCs, indicate that the critical behavior of SC Gd should belong to one of the short-range standard UCs. The MAPs, constructed using the CEs determined by NP, for the field range 1 to 2 T show significant but smaller deviation from quasi-linear variation. The MAPs for higher field-ranges show almost quasi-linear variation. This is the indication that the results obtained for the field range 1 to 2 T may be ambiguous. The critical analysis for above $T_{\rm{C}}$ have been ignored due to the reasons discussed in the main text while performing theoretical investigation on Ni system.
@@ 9. Scaling results
where $\tilde{h}(x)$ is the scaling function. Thus, the relation 11 states that for appropriate CEs and $T_{\rm{C}}$, the plots of $M(\mu_{0}H)^{1/\delta}$ vs. $\epsilon$$(\mu_{0}$)${}^{-\frac{1}{\beta+\gamma}}$ should correspond to a universal curve onto which all the isotherms collapse. The Eq. 11 may be formally converted as
@@
where $f_{+}$ and $f_{-}$ are defined for $T>T_{\rm{C}}$ and $T<T_{\rm{C}}$, respectively. The relation 12 states -- for appropriate values of $\beta$ and $\gamma$, the plots of renormalized magnetization, $m=M(H,\epsilon)|\epsilon|^{-\beta}$ vs. renormalized field, $h=\mu_{0}H|\epsilon|^{-(\beta+\gamma)}$ should correspond to collapse of the isotherms on two separate universal curves for below and above $T_{\rm{C}}$. Equation 12 is most commonly used relation to verify the appropriateness of the obtained CEs. It has been proven [17] that these scaling relations (Eqs. 11 and 12) can absorb significantly large errors appearing in the determined values of CEs. So, a new and appropriate relation is needed to check the reliability of determined CEs. However, Widom relation, $\delta=1+\displaystyle\frac{\gamma}{\beta}$, may be used as a primary tool to check the reliability and accuracy of determined CEs.
@@
Figures S8(a) to S8(d) show the plots of $m$ vs. $h$ constructed using the CEs determined by SP, IM, MIM and NP for Ni system. As one can see clearly that the $m$ vs. $h$ plots using the CEs determined by IM [Fig. S8(b)] show maximal deviation from overlapping on two separate isotherms. The $m$ vs. $h$ using the CEs determined by SP [Fig. S8(a)] have perfect collapse on two universal curves for the whole range of the M-H isotherms. Similar perfect collapse on two universal curves for the whole range of the M-H isotherms have been observed using the CEs determined by MIM [Fig. S8(c)] and NP [Fig. S8(d)] for below $T_{\rm{C}}$. We have used the same CEs, determined by MIM and NP for below $T_{\rm{C}}$, to construct the $m$ vs. $h$ plots for above $T_{\rm{C}}$, which show slight deviation from collapsing on a universal curve. Figures S8(e) to S8(h) are the renormalized Arrott-plots (RAPs) constructed using the CEs determined by SP, IM, MIM and NP. The perfect overlapping is closely observable for the CEs determined by SP, and MIM and NP for below $T_{\rm{C}}$. The RAPs constructed using the CEs determined by IM, MIM for above $T_{\rm{C}}$ and NP for above $T_{\rm{C}}$ show slightly large deviation from collapsing on universal curve. These are the indication that the CEs determined using SP, and MIM and NP for below $T_{\rm{C}}$ are appropriate to reveal the properties of a system under investigation. In fact, the CEs determined from the above three mentioned methods (SP, MIM and NP) match significantly to each other within error limit. However, no concrete conclusion can be made from the use of relation 11 because the constructed plots using the CEs determined by all methods overlap significantly onto a single universal curves as shown in Figs. S8(i) to S8(l). From these investigations, one important information is coming out -- the overlapping of all the isotherms on universal curves for entire field-range is the indication of the existence of single exchange coupling in the system, which was intentionally incorporated during simulation.
@@
Figure S9 represents the scaling plots and RAPs for SC Gd. One should note that the critical analysis on SC Gd have been carried out only for below $T_{\rm{C}}$ but we have plotted the above $T_{\rm{C}}$ isotherms also using the same CEs as determined for below $T_{\rm{C}}$. The plots shown in Fig. S9, have been constructed using the CEs determined by NP. The $m$ vs. $h$ plots show nice overlapping in the field-ranges which were used to determine CEs for below $T_{\rm{C}}$. However, in the low field range (approximately below 0.5 T) the $m$ vs. $h$ plots [Figs. S9(a) to S9(f)] show deviations from collapsing on the universal curves. In the similar fashion, the RAPs [Figs. S9(g) to S9(l)] also sow overlapping in the high field range while deviations are observable in the low field range. The scaling plots [Figs. S9(m) to S9(r)], constructed using the relation 11, also show significant deviations in the low-field range mainly below $T_{\rm{C}}$. The reported scaling results using Eq. 11 do not show deviation from collapsing on single universal curve. The reason behind this could be the skipping of the low-field data. However, from the deviation in the low-field region, we state that this is the signature of the competing energies along with symmetric exchange interaction in SC Gd.